OpenStudy (anonymous):

The school cafeteria sells two kinds of wraps: vegetarian and chicken. The vegetarian wrap costs $1.00 and the chicken wrap costs $1.80. Today they made $98.80 from the 70 wraps sold. How many of the wraps sold were vegetarian?

Parth (parthkohli):

\(x + 1.80y = 98.80\) \(x + y = 70\) \(\text{Do you know how to solve systems?} \)

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Can you show me how too

OpenStudy (anonymous):

the anwser is 34 vegeterian wraps. To solve the system express y from the second equation with x y=70-x. Substitute y in the second equation and you got it.

OpenStudy (anonymous):

oh ok i get it now